OpenStudy (anonymous):

Lisa will make punch that is 25% fruit juice by adding pure fruit juice to 2L mixture that is 10% pure fruit juice. How many liters of pure fruit juice does she need to add?

OpenStudy (mertsj):

.1(2)+x=.25(2+x)

OpenStudy (anonymous):

100p + 10* 2 = 25 ( p + 2)

OpenStudy (anonymous):

Can you solve from here?

OpenStudy (anonymous):

yes thanks